1. Considering the following reaction: 3Si + 2n2 →Si3N4

2. When 26.8 moles of Si react with 21.845 moles of N2, how many moles of Si3N4 are formed?

3. What is the limiting reactant?



1
Expert's answer
2021-11-05T09:32:07-0400

Mole ratio of Si:Si3N4= 3:1

= 26.8/3= 8.933

Mole ratio N2: Si3N4=2:1

= 21.845/2= 10.923

= 10.923+8.933= 19.856moles